OpenStudy (anonymous):

use trigonometric identities to solve the following equation for 0

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ 2(\sin θ)^2+\cos θ-1=0}\) \(\normalsize\color{blue}{ 2(\sin^2θ)+\cos θ-1=0}\) \(\normalsize\color{blue}{ 2(1-\cos^2θ)+\cos θ-1=0}\) \(\normalsize\color{blue}{ 2-2\cos^2θ+\cos θ-1=0}\) \(\normalsize\color{blue}{ -2\cos^2θ+\cos θ+1=0}\)

OpenStudy (solomonzelman):

Let cos(theta) = a, and solve

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ -2a^2+a+1=0}\) \(\normalsize\color{blue}{ 2a^2-a-1=0}\) \(\normalsize\color{blue}{ (2a+1)(a-1)=0}\) \(\normalsize\color{blue}{a=~-½ ,~~~~1}\)

OpenStudy (solomonzelman):

So, \(\normalsize\color{blue}{\cosθ=-1/2}\) or \(\normalsize\color{blue}{\cosθ=1}\)

OpenStudy (solomonzelman):

Let me know if you need more help.

OpenStudy (anonymous):

That's great. Thanks! I just needed the first one or two steps, but now I can compare and make sure I'm doing the rest of it correct. Thanks again.

OpenStudy (solomonzelman):

Anytime, for the first step, just in case.... \(\normalsize\color{black}{\sin^2θ+\cos^2θ=1~~~~~~~~\rm identity}\) and then you subtract cos²θ from both sides, to get that \(\normalsize\color{black}{\sin^2θ=1-\cos^2θ}\) and then I substituted for sin²θ to put everything in terms of cosines.

OpenStudy (solomonzelman):

have fun :)